TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes


Created by Main.MirkoPojer on 2007-06-04 - 18:44

Powering Tests Sector 78 - Afternoon Meeting 04/06/2007

Test Report Monday 04/06/2007

  • RQ5.L8: Powered up to nominal. During the first trial, the quench detector triggered at 27A. The magnet was powered at its nominal current for 4 hours. Tests successful. MPP is asked to give permission to increase the detection threshold to 100 mV (nominal) to carry out the Fast Power Abort test (PNO.9)

  • RQ4.L8:
    • Powered up to 250 A while RQ5.L8 was at nominal in order to check the integrity of the circuit. Test OK.
    • Once RQ5.L8 test was completed and with the agreement of MPP, RQ4.L8 was powered at different current levels up to its nominal current (250, 600, 2000 and 3610 A) and kept there for 20 minutes. Test OK.
    • Squeezing tests (PSQ). Test OK.
    • Fast Power Abort from PIC at nominal current (PNO.9). The circuit did not quench, however, AB/PO considers the test as failed due to the FWD circuit creating a fault. Under investigation. Test NOT OK.

  • RQ6.R7B1 and RQ6.R7B2
    • PCC Circuit Config tests going on. Some correction in the quench detectors have to be carried out. The noise level in the detector looks promising. Tests will continue tomorrow Tuesday.

  • RD2.R8
    • Amalia has analysed previous quenches in this magnet. None of them showed any effect on the level and temperatures in the DFBMA, hence they can conclude that the quench ocurred in the DFBMA but they cannot confirm its cause yet.
    • AB/CO has completed the interlock and circuit synoptics modifications needed for the new instrumentation in D2. Reiner will get the hardwire ready as soon as he finishes the correction in Q6.

  • Boil-off Tests
    • DFBAO Low Current Module: test completed, further studies need to be done in the He level gauges
    • DFBAO High Current Module: Test going out. They should be completed this evening. Promising results so far.
    • DFBAN being refilled, the calibration data is being checked.

Planning for Tuesday (05/06/2007)

  • RQ5.L8:
    • Squeezing Tests (PSQ)

  • RQ4.L8:
    • Run at nominal current for some hours
    • Repeat PNO.9 (FPA) if required by AB/PO

  • RQ4.L8 & RQ5.L8 & Correctors
    • Powering at nominal current together (?)

  • ElQA RB.A78
    • Tests will be restarted tomorrow Tuesday. First priority activity.

  • 60A Converter Type test
    • If possible in coordination with the ElQA.

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-06-04 - 11:22

Short Circuit Test in UA27 - 04.06.07

PDF RAT_20070604.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint2x2007x06x04x11x16/RAT_20070604.pdf

Microsoft Excel Spreadsheet SOC_24hr_heatrun_UA27.xls:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint2x2007x06x04x11x16/SOC_24hr_heatrun_UA27.xls

************************

Thursday 07.06: 24hr heat run - SOC attached.

************************

EE - IST    
13kA EE DONE 16-04-07
600A EE DONE 23-04-07
     
Circuits >2kA    
Powering tests PC 4-8kA DONE 11-05-07
8h run 4-8kA DONE 29-05-07
Interlock tests RQD&F DONE 30-05-07
Powering tests RQD&F DONE 30-05-07
Preparation RB.A23 DONE 19-04-07
Interlock tests RB.A23 DONE 25-05-07
Powering tests RB.A23 DONE 30-05-07
8h run >= 8kA ONGOING 04-06-07
     
Circuits < 2kA    
120A powering tests DONE 03-05-07
600ANOEE powering tests DONE 08-05-07
600A-EE interlock tests DONE 03-05-07
600A-EE powering tests DONE 18-05-07
8 hr run PC < 2kA DONE 29-05-07
     
24hr heat run test - 07-06-07
     
Polarity Check < 8KA - 05-06-07
AUG- UPS test - 08-06-07

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-06-04 - 11:07

Powering Tests Sector 78 - Minutes 04/06/2007

PDF RAT_20070604_2.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x06x04x11x06/RAT_20070604_2.pdf

Pptm D2_6000A.pptm:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x06x04x11x06/D2_6000A.pptm

Comments:

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2007-06-04 - 08:54

Cooldown Sector 78

The commissioning programme on Q4, Q5, D2 and the associated correctors continued during this week. While the Q4, Q5 and D2 commissioning was completed, the commissioning of the 120 A correctors associated to these magnets could not be completed: some additional tuning of the software in the FGC remains to be done.

Yesterday (Thursday May30th) Q4, Q5, D2 were taken together to nominal current. At about 80% of the ramp, D2 quenched (5268 A). As expected, this caused the shutdown of the converters in the matching section and during the ramp down Q4 and Q5 also quenched. The data of this event is still being analysed.

The Electrical Quality Assurance team qualified the focusing and the defocusing quadrupole circuits and the matching section quadrupole Q6 at Point 7.

All the DFBs of the sector have been equipped with dry-air bags to ensure no condensation forms on the current leads and their chimneys.

Unfortunately, on Friday (June 1st) morning, the safety chain of cabinet was accidentally broken and the sector was disconnected from the pumping group. The cryo team is now busy recovering the situation to re-establish nominal temperature on the magnets for Monday morning.

Comments:

Edit, attach file or add comment


Created by Main.UnknownUser on 2007-06-04 - 00:37

RAT in P4 - Minutes 1st June 2007 (3 weeks for Cool-Down !)

PDF RAT_20070601.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x06x04x00x35/RAT_20070601.pdf

Comments:

Edit, attach file or add comment

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2007-06-01 - JacekSzkutnik
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login